OpenStudy (anonymous):

Find an exact solution for Square root of 5 - 1/x = square root of 5/2 Then find the approximate solution to the nearest tenth.

OpenStudy (anonymous):

\[\frac{ \sqrt{5}-1 }{ x } = \frac{ \sqrt{5} }{ 2 }\]

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

How do you cross multiply with \[\sqrt{5}\] 's?